OpenStudy (anonymous):

can anyone help me solve this: f(x)= 1/3x^2-2x f(x+h)=?

OpenStudy (amistre64):

just replace 'x' with 'x+h'

OpenStudy (amistre64):

1/3x^2 -2x 1/3(x+h)^2 -2(x+h)

OpenStudy (anonymous):

f(x+h)=1/3*(x+h)^2-2(x+h)

OpenStudy (amistre64):

how the equation looks I cant tell cause of the way its presented

OpenStudy (amistre64):

\[f(x)=\frac{1}{3x^2 -2x}\] \[f(x+h)=\frac{1}{3(x+h)^2 -2(x+h)}\]

OpenStudy (anonymous):

amistre, how do u get those horizontal lines for division symbol ?

OpenStudy (anonymous):

magic!

OpenStudy (amistre64):

i think about it real hard like lol

OpenStudy (amistre64):

\[\frac{top}{bottom}\_] ^ remove the '_'

OpenStudy (amistre64):

And \({this \over that}\_) works for simple stuff

OpenStudy (amistre64):

We can see that \(this\over that\) is for inline typing and \[\frac{top}{bottom}\] breaks the line

OpenStudy (anonymous):

how to evaluate this limit g(x)= √2x+1 lim (g(x+h)-g(x))/h h---0

OpenStudy (anonymous):

what is √2x+1